The length of a regular pool is 6 meters less than twice its width. If the pool's perimeter is 126 meters, what are its dimensions?
Answer by rapaljer(4671)

Let x = width of the pool
2x-6 = length of the pool

2W + 2L = Perimeter
2(x) + 2(2x-6) = 126

2x + 4x - 12 = 126
6x - 12 = 126
6x=138
x=138/6
x=23 meters WIDTH
2x-6=46-6=40 meters LENGTH

Chcck: 2*23 + 2*40=46 + 80 = 126 It checks!!!

For additional explanation on WORD PROBLEMS, please see my own website. The easiest way to find it is to use the easy to spell and easy to remember link www.mathinlivingcolor.com. On this single page website, there is a link at the bottom of the page that takes you to my Homepage.

Near the top of my homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time", and choose Basic Algebra (Chapter 1). There you will find explanations that may be easier to understand than your traditional textbooks, and many of the exercises from these sections are solved in color on the "MATH IN LIVING COLOR" pages that go with these sections.
If you like my website, please tell your friends and family. It's all FREE. You can contact me by Email at rapaljer@seminolestate.edu.

Dr. Robert J. Rapalje, Retired
Professor of Mathematics
Seminole State College of Florida